# Deep flows transmitted by forced surface gravity waves

### v1

@@ Abstract
We examine a deep-water surface gravity wave packet generated by a pressure disturbance in the Lagrangian reference frame. The pressure disturbance has the form of a narrow-banded weakly nonlinear deep-water wave packet. During the forcing, the vorticity equation implies that the momentum is entirely in the near-surface Lagrangian mean flow. After the forcing is turned off, the wave packet propagates away from the forcing region carrying with it the energy imparted from the pressure forcing. These waves together with their induced long wave response have no momentum, in agreement with the classical results of Longuet-Higgins & Stewart (1962) and McIntyre (1981). The momentum imparted to the fluid from the form stress of the pressure forcing is entirely in a dipolar structure that remains in the area of the forcing event. This is in contrast with the finite-depth scenario discussed in McIntyre (1981) where the momentum imparted to the fluid by the forcing event propagates away as shallow water waves. Finally, we examine waves propagating from deep to shallow water and argue that wave packets, which initially have no momentum, may have non-zero momentum in finite-depth water through the trapping of a long wave at the step change in depth, which ensures the total momentum of the system remains conserved.
@@ 1. Introduction
Much of the ocean’s momentum is transferred from the atmosphere by form stresses pressing on the faces of surface gravity waves (Grare et al., 2013; Melville, 1996). Yet the fluid dynamical details of how atmospheric pressure is transmitted across the surface — somehow simultaneously forcing a current beneath the surface while also exciting a surface gravity wave — are mysterious. In this paper we elucidate some of these details through a detailed asymptotic analysis of the forced growth of a surface gravity wave packet of finite extent by an idealized atmospheric pressure distribution.
@@
The unforced cousin of this problem was analyzed by Longuet-Higgins & Stewart (1962), who computed the total second-order mean flow associated with a freely-propagating deep-water surface gravity wave packet. Longuet-Higgins & Stewart (1962) showed that the surface-wave-induced mean flow has two components: first, a shallow surface component with no Eulerian-mean signature wholly attributable to the surface gravity wave Stokes drift, which may be evaluated diagnostically from and has similar vertical scales as the primary wave field. The second component of the mean flow is driven by the convergence and divergence of the primary flow, is purely Eulerian-mean, and is longer and deeper with scales similar to that of the wave packet. Remarkably and as emphasized by McIntyre (1981), the total vertically-integrated momentum of the freely-propagating packet vanishes at every horizontal point. In other words, the momentum carried by the deep return flow exactly cancels that of the shallow Stokes drift — implying that deep-water surface gravity waves do not carry momentum with them as they propagate.
@@
The absence of momentum in freely-propagating wave packets begs the question: where is the momentum transferred from the atmosphere during wave generation? McIntyre (1981) argues that the momentum transferred during wave generation is carried away by packet-scale, shallow surface gravity waves — which being shallow, can carry momentum. But McIntyre (1981)’s answer, in addition to lacking the details that might accompany a more systematic solution of the equations of motion, produces counter-intuitive outcomes if the depth of the fluid is increased while holding the scale of the packet and atmospheric forcing fixed. For example, we find that atmospheric momentum would have to be instantaneously transferred farther and farther from the disturbance, carried by shallow water waves with longer and longer wavelengths without bound as the depth of the fluid increases.
@@
This paper revisits this question through a detailed and systematic asymptotic solution of the Lagrangian equations of motion. Our aim is to elucidate the fate of atmospheric momentum transferred to a fluid during surface gravity wave packet generation, and in process, clarify McIntyre (1981)’s argument. We find, contrary to McIntyre (1981), that in an infinitely deep, non-rotating, and unstratified fluid, atmospheric momentum is not transferred to propagating waves but instead remains in place after forcing ceases while the packet propagates away. Moreover, we find the precise mechanism by which this occurs: as the packet grows, the pressure forces that drive its deep return flow are opposed — pointwise within the interior of the fluid — by the packet-scale component of the atmospheric pressure gradient. As a result, there is no deep return flow while the packet is forced, and it is only at the moment the forcing ceases that the deep return flow and its counterpart — the part that contains the net momentum transferred from the atmosphere — remains behind. We call this part of the flow, which is effectively transmitted from the surface into the deep ocean during the generation of surface gravity waves, and is given precisely as the negative of the deep return flow associated with the propagating packet at the moment the forcing ceases, as the “wave-transmitted flow”.
@@
An additional point of confusion arises from a close reading of Longuet-Higgins & Stewart (1962) and McIntyre (1981). Our analysis shows that when both the primary wave packet and the long wave response are in deep-water, neither carries momentum. Yet, shallow water waves can carry momentum, which seemingly leads to a violation of conservation of momentum if the wave packet propagates from deep to shallow water. To better understand this, we consider a wave packet in infinitely deep-water which experiences a step to finite depth and argue that this leads to the long wave being partially trapped at the step. When the momentum in each component is accounted for, the total momentum of the system remains conserved.
@@
The outline of this paper is as follows. In §2 we present the equations of motion and several conservation laws for deep-water weakly nonlinear narrow banded surface gravity waves in the Lagrangian reference frame. In §3, we satisfy the equations of motion to third order in a small parameter. §4 considers the evolution of the waves and the mean flow after the forcing is turned off. The conclusions are presented in §6.
@@ 2. Equations of motion
where $(x,y)$ are the particle locations, $(a,b)$ are particle labels, $p$ is the fluid pressure, $g$ is the acceleration due to gravity, and $\mathcal{J}$ is defined as
@@
Derivatives with respect to the time $\tau$ take $(a,b)$ to be fixed, so that $\partial_{\tau}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}D/Dt$. Even though the acceleration term is linear in the Lagrangian frame, the pressure term is now nonlinear. This highlights the fact that the mapping between the two frames may be (highly) nonlinear.
@@
In the absence of forcing, we require that the pressure $p$ vanish at the free surface, i.e. $p=0$ at $b=0$. Finally, we require that $y_{\tau}\to 0$ as $b\to-\infty$ so that there is no flow through the infinitely deep bottom.
@@ 2.1. Conservation laws
If we vertically integrate equation (1), we have
(12) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b+\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b=0.$
We can rewrite the terms involving the pressure by noting
(13) $\frac{\partial}{\partial a}\frac{1}{2}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\frac{1}{2}\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}+py_{ab}-p_{ab}y\right){\rm d}b.$
Integration by parts implies
(14) $\frac{1}{2}\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b+\frac{1}{2}\left(py_{a}-yp_{a})\right|_{b=0},$
so that the conservation of horizontal momentum becomes
(15) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}\ {\rm d}b+\frac{\partial}{\partial a}\int_{-\infty}^{0}py_{b}{\rm d}b=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0}.$
In the absence of pressure forcing at the surface, $p=p_{a}=0$ at the free surface so that the right hand side becomes zero. The second term here represents the momentum flux in this framework, which we define as $S$, where
@@
(17) $\frac{\partial\mathcal{I}}{\partial\tau}+\frac{\partial S}{\partial a}=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0},$
with $\mathcal{I}$ defined as
@@
(19) $\frac{\partial\mathcal{I}}{\partial\tau}+\frac{\partial S}{\partial a}=0,$
@@
We derive an energy integral by multiplying equation (1) by $x_{\tau}$ and equation (2) by $y_{\tau}$ and vertically integrating to find
(20) $\frac{\partial}{\partial\tau}\int_{-\infty}^{0}\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\rm d}b+\int_{-\infty}^{0}\left(p_{a}(x_{\tau}y_{b}-y_{\tau}x_{b})+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right){\rm d}b=0.$
@@
(21) $\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y)\right){\rm d}b.$
@@
(22) $\begin{split}&\int_{-\infty}^{0}\left(p_{a}(x_{\tau}y_{b}-y_{\tau}x_{b})+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right){\rm d}b\\ &\qquad=\frac{\partial}{\partial a}\int_{-\infty}^{0}\left(p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y)\right){\rm d}b+\left.\left(p_{a}(xy_{\tau}-x_{\tau}y)\right)\right|_{b=0},\end{split}$
@@
(23) $\frac{\partial E}{\partial\tau}+\frac{\partial F}{\partial a}=\left.\left(p_{a}(x_{\tau}y-xy_{\tau})\right)\right|_{b=0},$
@@
Next, consider the kinematic relationship $z_{\tau}^{\star}=\chi_{z}$, for $\star$ denoting complex conjugate and $\chi=\phi+\mathrm{i}\psi$ the complex velocity potential. If we then map to the Lagrangian frame, and let $s=a+\mathrm{i}b$, we have, $z_{\tau}^{\star}z_{w}=\chi_{w}$ (Longuet-Higgins, 1980b) so that the stream function $\psi$ in the Lagrangian reference frame satisfies
@@ 2.2. Momentum and energy transfer by surface pressure forcing
(28) $\frac{\partial}{\partial\tau}\iint\mathcal{J}u{\,\rm d}a{\,\rm d}b=\int py_{a}\,|_{b=0}{\,\rm d}a\,.$
@@
(29) $\frac{\partial}{\partial\tau}\iint\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\,\rm d}a{\,\rm d}b=\int\left.p(x_{\tau}y_{a}-x_{a}y_{\tau})\right|_{b=0}{\rm d}a.$
@@ 3. Weakly nonlinear narrow-banded waves
In our analysis we take $\mathcal{J}=1$, which amounts to a gauge choice (Salmon, 2020) sometimes known as Miche’s condition (Clamond, 2007).
The pressure forcing $p^{\mathrm{atm}}$ is defined at the free surface, $b=0$, at first order to be
@@
and $\mathcal{A}=\mathcal{A}(\epsilon a,\epsilon\tau)$ is a slowly varying function (with units of length), which represents the modulation to the forcing event in space and time.
@@ 3.3. Second order expansion
Unlike in the Eulerian frame, at second order these equations have no wavy terms and therefore represent the mean flow response to the pressure forcing.
@@
The mean momentum of the system is
@@ 4. Behavior after forcing is turned off
We now turn off the forcing at a time $\tau=T$, allowing us to follow Salmon (2020) closely. At first order,
@@
which implies we can define a stream function for the second order mean flow:
@@
(61) $\Delta\psi=-2k^{2}\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb}.$
@@
We can consider the particular and homogeneous solution to this system. That is, let $\psi=\psi_{c}+\psi_{LW}$ where $c$ denotes carrier and $LW$ the long wave contribution to the stream function. By inspection,
(65) $\psi_{c}=-\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb},$
agreeing with the classical prediction of Stokes. This implies $\psi_{LW}$ must obey the relationship
(66) $\Delta\psi_{LW}=0$
with boundary condition $\psi_{LW}(b=0)=-4\omega^{3}T^{2}\mathcal{A}^{2}$. The domain of the labels here is the lower half plane, so that we may use the theory of Poisson kernels to write down the solution
(67) $\psi_{LW}=-\frac{4\omega^{3}T^{2}k^{2}}{\pi}\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}b}{(a^{\prime}-a)^{2}+b^{2}}{\rm d}a^{\prime}.$
$\psi_{c}$ and $\psi_{LW}$ are used to generate the figures in this manuscript. Note, some care must be taken when considering these functions at $b=0$, as $\psi_{LW}$ is not zero there, but instead reduces to the prescribed boundary condition with $b(1+(a-a^{\prime})^{2}/b^{2})^{-1}$ being a $\delta-$function sequence as $b\to 0$.

### v2

@@ Abstract
We examine a two-dimensional deep-water surface gravity wave packet generated by a pressure disturbance in the Lagrangian reference frame. The pressure disturbance has the form of a narrow-banded weakly nonlinear deep-water wave packet. During forcing, the vorticity equation implies that the momentum resides entirely in the near-surface Lagrangian-mean flow, which in this context is often called the “Stokes drift”. After the forcing turns off, the wave packet propagates away from the forcing region, carrying with it most of the energy imparted by the forcing. These waves together with their induced long wave response have no momentum in a depth integrated sense, in agreement with the classical results of Longuet-Higgins & Stewart (1962) and McIntyre (1981). The total flow associated with the propagating packet has no net momentum, in agreement with the classical results of Longuet-Higgins & Stewart (1962) and McIntyre (1981). In contrast with the finite-depth scenario discussed by McIntyre (1981), however, momentum imparted to the fluid during forcing resides in a dipolar structure that persists in the forcing region — rather than being carried away by shallow water waves. We conclude by examining waves propagating from deep to shallow water and show that wave packets, which initially have no momentum, may have non-zero momentum in finite-depth water through reflected and trapped long waves. This explains how deep water waves acquire momentum as they approach shore. The artificial form of the parameterized forcing from the wind facilitates the thought experiments considered in this paper, as opposed to striving to model more realistic wind forcing scenarios.
@@ 1. Introduction
Much of the ocean’s momentum is transferred from the atmosphere by form stresses pressing on the faces of surface gravity waves (Grare et al., 2013; Melville, 1996). Yet the fluid dynamical details of how atmospheric pressure is transmitted across the surface — somehow simultaneously forcing a current beneath the surface while also exciting a surface gravity wave — are mysterious (Pizzo et al., 2021). In this paper we elucidate some of these details through a detailed asymptotic analysis of the forced growth of a two-dimensional surface gravity wave packet of finite extent by an idealized atmospheric pressure distribution.
@@
The unforced cousin of this problem was analyzed by Longuet-Higgins & Stewart (1962), who computed the total second-order mean flow associated with a freely-propagating deep-water surface gravity wave packet. Longuet-Higgins & Stewart (1962) showed that the surface-wave-induced mean flow has two components: first, a shallow surface component with no Eulerian-mean signature wholly attributable to the surface gravity wave Stokes drift, which may be evaluated diagnostically from and has similar vertical scales as the primary wave field. The second component of the mean flow is driven by the convergence and divergence of the primary flow, is purely Eulerian-mean, and is longer and deeper with scales similar to that of the wave packet. Remarkably and as emphasized by McIntyre (1981), the total vertically-integrated momentum of the freely-propagating packet vanishes at every horizontal point.
@@
The absence of momentum in freely-propagating wave packets begs the question: where is the momentum transferred from the atmosphere during wave generation? McIntyre (1981) argues that the momentum transferred during wave generation is carried away by packet-scale, shallow surface gravity waves — which being shallow, can carry momentum. But McIntyre (1981)’s answer, in addition to lacking the details that might accompany a more systematic solution of the equations of motion, produces counter-intuitive outcomes if the depth of the fluid is increased while holding the scale of the packet and atmospheric forcing fixed. For example, in McIntyre (1981)’s theory atmospheric momentum would have to be instantaneously transferred farther and farther from the disturbance, carried by shallow-water waves with longer and longer wavelengths without bound as the depth of the fluid increases.
@@
This paper revisits the momentum question through a detailed and systematic asymptotic solution of the Lagrangian equations of motion. Our aim is to elucidate the fate of atmospheric momentum transferred to a fluid during surface gravity wave packet generation, and in process, clarify McIntyre (1981)’s argument. We find, contrary to McIntyre (1981), that in an infinitely deep, non-rotating, and unstratified fluid, atmospheric momentum is not transferred to propagating waves but instead remains in place after forcing ceases while the packet propagates away. Moreover, we find the precise mechanism by which this occurs: as the packet grows, the pressure forces that drive its deep return flow are opposed — pointwise within the interior of the fluid — by the packet-scale component of the atmospheric pressure gradient. As a result, there is no deep return flow while the packet is forced, and it is only at the moment the forcing ceases that the deep return flow and its counterpart — the part that contains the net momentum transferred from the atmosphere — remains behind. We call this part of the flow, which is effectively transmitted from the surface into the deep ocean during the generation of surface gravity waves, and is given precisely as the negative of the deep return flow associated with the propagating packet at the moment the forcing ceases, the “wave-transmitted flow”.
@@
An additional point of confusion arises from a close reading of Longuet-Higgins & Stewart (1962) and McIntyre (1981). Our analysis shows that when both the primary wave packet and the long wave response are in deep-water, neither carries momentum. Yet, shallow-water waves can carry momentum, which seemingly leads to a violation of conservation of momentum if the wave packet propagates from deep to shallow-water. To better understand this, we consider a wave packet in infinitely deep-water which experiences a step to finite depth and argue that this leads to the long wave being partially trapped at the step. When the momentum in each component is accounted for, the total momentum of the system remains conserved.
@@
The outline of this paper is as follows: in §2 we present the equations of motion and several conservation laws for deep-water weakly nonlinear narrow banded surface gravity waves in the Lagrangian reference frame. In §3, we satisfy the equations of motion to third order in a small parameter. §4 considers the evolution of the waves and the mean flow after the forcing is turned off. The conclusions are presented in §5.
@@ 2. Equations of motion
where $(x,y)$ are the particle locations, $(a,b)$ are particle labels, $p$ is the fluid pressure, $g$ is the acceleration due to gravity, and the specific volume $\mathcal{J}$ is defined as
@@
Derivatives with respect to the time $\tau$ take $(a,b)$ to be fixed, so that $\partial_{\tau}\stackrel{{\scriptstyle\mathrm{def}}}{{=}}D/Dt$. Even though the acceleration term is linear in the Lagrangian frame, the pressure gradient terms in (1)–(2) are now nonlinear. This highlights the fact that the mapping between the two frames may be (highly) nonlinear.
@@
In the absence of forcing, we require that the pressure $p$ vanish at the free surface, i.e. $p=0$ at $b=0$. Finally, we require that $y_{\tau}\to 0$ as $b\to-\infty$ so that flow vanishes at depth. A word of warning concerning nomenclature – we say waves are deep when $kd\gg 1$ for $k$ their wavenumber and $d$ the depth of the fluid while we say a wave packet is in deep-water if $\Delta kd\gg 1$ for $\Delta$ the non-dimensional bandwidth of the wave packet, here taken to be $\epsilon$. There can be scenarios where the waves themselves are in deep water, but the wave packets are not.
@@ 2.1. Conservation laws
Vertically integrating equation (1) yields
(12) $\partial_{\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}{\rm d}b+\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b=0.$
We rewrite the second term in (12) by noting
(13) $\frac{1}{2}\partial_{a}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\frac{1}{2}\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}+py_{ab}-p_{ab}y\right){\rm d}b.$
Further integrating by parts produces
(14) $\frac{1}{2}\partial_{a}\int_{-\infty}^{0}\left(py_{b}-p_{b}y\right){\rm d}b=\int_{-\infty}^{0}\left(p_{a}y_{b}-p_{b}y_{a}\right){\rm d}b+\frac{1}{2}\left(py_{a}-yp_{a})\right|_{b=0},$
so that the $b$-integrated conservation of horizontal momentum expressed by equation (12) becomes
(15) $\partial_{\tau}\int_{-\infty}^{0}\mathcal{J}x_{\tau}\ {\rm d}b+\partial_{a}\int_{-\infty}^{0}py_{b}{\rm d}b=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0}.$
In the absence of pressure forcing at the surface, $p=p_{a}=0$ at the free surface so that the right hand side becomes zero. The second term here represents the $x$-momentum flux in this framework, which we define as $S$, where
@@
(17) $\mathcal{I}_{\tau}+S_{a}=\frac{1}{2}\left(yp_{a}-py_{a})\right|_{b=0},$
with the impulse $\mathcal{I}$ defined as
@@
(19) $\mathcal{I}_{\tau}+S_{a}=0$
@@
[…]
(20) $0=\partial_{\tau}\int_{-\infty}^{0}\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\rm d}b+\int_{-\infty}^{0}\left[p_{a}\left(x_{\tau}y_{b}-y_{\tau}x_{b}\right)+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right]\,{\rm d}b\,.$
@@
(21) $\partial_{a}\int_{-\infty}^{0}p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y){\rm d}b\,.$
@@
(22) $\begin{split}&\int_{-\infty}^{0}\left[p_{a}(x_{\tau}y_{b}-y_{\tau}x_{b})+p_{b}(y_{\tau}x_{a}-x_{\tau}y_{a})\right]\,{\rm d}b\\ &\qquad=\partial_{a}\int_{-\infty}^{0}\left[p_{b}(xy_{\tau}-x_{\tau}y)+p(xy_{\tau b}-x_{\tau b}y)\right]\,{\rm d}b+\left[p_{a}\left(xy_{\tau}-x_{\tau}y\right)\right]_{b=0},\end{split}$
@@
(23) $E_{\tau}+F_{a}=\left[p_{a}\left(x_{\tau}y-xy_{\tau}\right)\right]_{b=0},$
@@
Next, consider the kinematic relationship $z_{\tau}^{\star}=\chi_{z}$, for $\star$ denoting complex conjugate and $\chi=\phi+\mathrm{i}\psi$ the complex velocity potential. If we then map to the Lagrangian frame, and let $s=a+\mathrm{i}b$, we have, $z_{\tau}^{\star}z_{s}=\chi_{s}$ (Longuet-Higgins, 1980b) so that the stream function $\psi$ in the Lagrangian reference frame satisfies
@@ 2.2. Momentum and energy transfer by surface pressure forcing
(28) $\partial_{\tau}\iint\mathcal{J}u{\,\rm d}a{\,\rm d}b=\int py_{a}\,|_{b=0}{\,\rm d}a\,.$
@@
(29) $\partial_{\tau}\iint\mathcal{J}\left(\frac{x_{\tau}^{2}+y_{\tau}^{2}}{2}+y\right){\,\rm d}a{\,\rm d}b=\int\left.p(x_{\tau}y_{a}-x_{a}y_{\tau})\right|_{b=0}\,{\rm d}a\,.$
@@ 3. Weakly nonlinear narrow-banded waves
In our analysis we take $\mathcal{J}=1$, which amounts to a gauge choice (Salmon, 2020) sometimes known as Miche’s condition (Clamond, 2007). The pressure forcing $p^{\mathrm{atm}}$ is defined at the free surface, $b=0$, at first order to be
@@
and $\mathcal{A}=\mathcal{A}(\epsilon a,\epsilon\tau)$ is a slowly varying function (with units of length), which represents the modulation to the forcing event in space and time. For example, figure 3 is made with $\mathcal{A}=\exp\left\{-\left(a-c_{g}t\right)^{2}/2\delta^{2}\right\}$, where $c_{g}$ is the group velocity and the packet width $\delta$ is $O(\epsilon^{-1})$.
@@ 3.3. Second order expansion
Unlike in the Eulerian frame, at second order these equations have no wavy terms (in other words, no dependence on the rapidly-varying phase $\theta$) and therefore represent the mean flow response to the pressure forcing.
@@
The mean $x$-momentum of the system is
@@ 4. Behavior after forcing is turned off
We now turn off the forcing at a time $\tau=T$ and consider a time when the packet as propagate sufficiently far that we can divide the total flow into two components: a component $\psi_{C}$ that propagates (is “carried” by) with the packet, and a second component “wave-transmitted” component $\psi_{WT}$ that remains at the forcing site.
4.1. The mean flow component co-propagating with the wave packet
To compute $\psi_{C}$, we consider the mean flow associated with a freely-propagating packet, following Salmon (2020) closely. At first order,
@@ 4.1. The mean flow component co-propagating with the wave packet
which implies we can define a stream function for the second order mean flow (a linearization of equation 26):
@@
(61) $\Delta\psi_{C}=-2k^{2}\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb}.$
@@
We can consider the particular and homogeneous solution to this system. By inspection,
(65) $\psi_{S}=-\omega^{3}T^{2}\mathcal{A}^{2}e^{2kb},$
agreeing with the classical prediction of Stokes. Using $\psi_{C}=\psi_{S}+\psi_{D}$ and plugging (65) into (61) implies that $\psi_{D}$ obeys
(66) $\Delta\psi_{D}=0$
with boundary condition $\psi_{D}(b=0)=-4\omega^{3}T^{2}\mathcal{A}^{2}$. The domain of the labels here is the lower half plane, so that we may use the theory of Poisson kernels to write down the solution
(67) $\psi_{D}=-\frac{4\omega^{3}T^{2}k^{2}}{\pi}\int_{-\infty}^{\infty}\frac{\mathcal{A}^{2}b}{(a^{\prime}-a)^{2}+b^{2}}{\rm d}a^{\prime}.$
$\psi_{S}$ and $\psi_{D}$ are used to generate the figures in this manuscript. Note, some care must be taken at $b=0$, where $\psi_{D}$ is not zero but rather reduces to the prescribed boundary condition with $b(1+(a-a^{\prime})^{2}/b^{2})^{-1}$ being a $\delta-$function sequence as $b\to 0$.